OpenStudy (anonymous):

the graph of f(x)=1/x^2-c has a vertical asymptote at x=3. find c

jimthompson5910 (jim_thompson5910):

the vertical asymptote occurs when x^2 - c is 0 since x = 3 is the vertical asymptote, we can plug this in to solve for c x^2 - c = 0 3^2 - c = 0 c = ??

OpenStudy (anonymous):

c= 3^2?

jimthompson5910 (jim_thompson5910):

or c = 9, the simplified form

OpenStudy (anonymous):

thanks so much

jimthompson5910 (jim_thompson5910):

np